A longboarder (45.0 kg) is travelling on her 2.00 kg skateboard at 5.35 m/s. She jumps off the board, continuing in the same direction at 4.25 m/s. What is the velocity of the skateboard after she jumps off?
1
Expert's answer
2020-11-03T10:28:05-0500

The initial momentum of the system was

"p_i =m_Sv_i + m_lv_i = (m_s+m_l)v_i"

where "m_s = 2kg" is the mass of the skateboard, "m_l = 45kg" is the mass of the longboarder and "v_i = 5.35m\/s" is their initial speed. Thus, we have:


"p_i = (45kg+2kg)\times 5.35m/s = 251.45\space kg\cdot m/s"

After the jump the longboarder took away the following momentum:


"p_{l}' = m_lv_l'"

where "v_l' = 4.25m\/s" is the speed of the longboarder after the jump. Thus, obtain:


"p_{l}' = m_lv_l' = 45kg\times 4.25m/s = 191.25\space kg\cdot m/s"

Thus, the momentum of the skateboard after the jump is:


"p_s' = p_i - p_l' = 251.45\space kg\cdot m/s-191.25\space kg\cdot m/s = 60.2\space kg\cdot m/s"

The speed of the skateboard then is:


"v_s' = \dfrac{p_s'}{m_s} = \dfrac{60.2\space kg\cdot m/s}{2kg} = 30.1m/s"

Answer. 30.1 m/s.